simplymariex:

What key features of a quadratic graph can be identified and how are the graphs affected when constants or coefficients are added to the parent quadratic equations? Compare the translations to the graph of linear function. Create examples of your own to explain the differences and similarities.

mathmale:

You ask several questions at once here. Could you possibly focus on one thing at a time? Example: " how are the graphs affected when constants are added to the parent quadratic equations?" For example: How is the graph of y=x^2 affected when this equation is modified and now appears as y=x^2+2? How is the graph of y=x^2 modified and now appears as y=(x-3)^2?

mathmale:

You ask several questions at once here. Could you possibly focus on one thing at a time? Example: " how are the graphs affected when constants are added to the parent quadratic equations?" For example: How is the graph of y=x^2 affected when this equation is modified and now appears as y=x^2+2? How is the graph of y=x^2 modified if the function now appears as y=(x-3)^2?

simplymariex:

I don't mean to... It's how my question is set up by my teacher...

simplymariex:

But thank you.

simplymariex:

I'll close this and do one by one.

mathmale:

Thank you. I think you'll learn more that way.